OpenStudy (rositaluv):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 Help me? PLEASE!

jimthompson5910 (jim_thompson5910):

Let D = b^2 − 4ac This is known as the discriminant

jimthompson5910 (jim_thompson5910):

If the discriminant D is a perfect square, then ax^2 + bx + c can be factored

OpenStudy (rositaluv):

Can you explain the factoring process to me? I am having trouble?

jimthompson5910 (jim_thompson5910):

Compare 2x^2 + 7x + 3 to ax^2 + bx + c

OpenStudy (rositaluv):

I wouldn't use quadratic formula would I?

jimthompson5910 (jim_thompson5910):

what are the values of 'a', 'b' and 'c'?

jimthompson5910 (jim_thompson5910):

You can if you want

OpenStudy (rositaluv):

a would be 2x^2? I am completely clueless

jimthompson5910 (jim_thompson5910):

a = 2, b = 7 and c = 3

OpenStudy (rositaluv):

ahh ok so know I have 2x^2 +7x+3 would do I do with that?

jimthompson5910 (jim_thompson5910):

So D = b^2 - 4ac D = 7^2 - 4(2)(3) ... Plug in a = 2, b = 7 and c = 3 D = 49 - 24 D = 25

jimthompson5910 (jim_thompson5910):

is 25 a perfect square?

OpenStudy (rositaluv):

Yes

jimthompson5910 (jim_thompson5910):

so 2x^2 + 7x + 3 can be factored

OpenStudy (rositaluv):

And the ending value is?

jimthompson5910 (jim_thompson5910):

25

jimthompson5910 (jim_thompson5910):

they don't want you to factor, they just want to know if it's possible to factor or not

jimthompson5910 (jim_thompson5910):

they just want the value of b^2 - 4ac

OpenStudy (rositaluv):

I understand how you got the answer it is just they give me a blank to type the answer in and I'm not sure what the correct thing to put in it is.

jimthompson5910 (jim_thompson5910):

you type in 25

jimthompson5910 (jim_thompson5910):

that's the final value of b^2 - 4ac

jimthompson5910 (jim_thompson5910):

where a = 2, b = 7 and c = 3

OpenStudy (rositaluv):

ok thank god for your help you are a lifesaver! :D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (rositaluv):

Got a 100 thanks to you! I won't fail Algebra! tehe;)

jimthompson5910 (jim_thompson5910):

well you did most of the work, I only helped out with 1 problem

jimthompson5910 (jim_thompson5910):

congrats on getting a perfect score

OpenStudy (rositaluv):

Yeah I know its just thats like the first 100 in a while thanks Again! :)

jimthompson5910 (jim_thompson5910):

sure thing